1/3x-(1/2x-2)= 3 I am stumped on this one... how does the fraction multiply in -2.
I end up getting -1/6-2/3= 3 then 1/9= 3 but the answer is -6 and we are solving for x

Answer by Boreal(15235)

1/3x-(1/2x-2)= 3
(1/3)x-(1/2)x+2=3 distributing the - sign over the parentheses
(1/3)x-(1/2)x=1, subtracting 2.
(1/3)x=(2/6)x
(1/2)x=(3/6)x
Therefore, we have -(1/6)x=1
multiply both sides by 6
-x=6
x=-6